OpenStudy (anonymous):

what is 8-2(times)[(5-3)^(2) -1]= please explain all steps!!!????

OpenStudy (anonymous):

\[ 8 - 2\cdot[(5-3)^2-1] = 8 - 2 \cdot [2^2 - 1] = 8 - 2 \cdot 3 = 8 - 6 = \fbox{2}. \]